OpenStudy (anonymous):

Can i have some help with this I have the format i'm just not sure were to go from here. y=mx+b y =x3 -10 18. Write an equation of a line whose graph is parallel to the graph of y = 3x – 10

OpenStudy (anonymous):

To write a graph that is parallel, just write the same thing with a different y intercept but the same slope. For instance, y = 3x - 10 has a Slope of 3 and the y intercept is at -10. If you write another equation like this and just change the y intercept, it will be parallel.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

How do i change the y intercept?

OpenStudy (anonymous):

Just change the number

OpenStudy (anonymous):

To any number?

OpenStudy (anonymous):

Yes, to any real number

OpenStudy (anonymous):

It can be 1/2 or 2 or 55 or 2/3 just change the number

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So one possible solution is y = 3x + 2/3 OR y = 3x - 4

OpenStudy (anonymous):

is -4 a real number?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Kay so y= 3x -10 y=3x -4 y=14

OpenStudy (anonymous):

What are you trying to do? The question is Write an equation of a line whose graph is parallel to the graph of y = 3x – 10 The one answer is y = 3x-4

OpenStudy (anonymous):

I thought i was doing what you said. But i guess i wasn't idk

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So what do we do with y=+3 -14

OpenStudy (anonymous):

That is your answer to the question

OpenStudy (anonymous):

See here https://www.desmos.com/calculator/fdestjraio

OpenStudy (anonymous):

Oh okay i meant to write 10 though XD

OpenStudy (anonymous):

Thanks for your time and the link. ^ ^

OpenStudy (anonymous):

yw